OpenStudy (anonymous):

Chad has a rope that is 20 yards long. How many pieces of rope measuring 4/7 of a yard can he divide his rope into?

OpenStudy (anonymous):

35

OpenStudy (texaschic101):

4/7x = 20 x = 20/(4/7) x = 20 * 7/4 x = 140/4 x = 35